Find a solution u(x,t) to the problem ๐œ•๐‘ข ๐œ•๐‘ก = 1.71 ๐œ•^2๐‘ข/๐œ•๐‘ฅ^2 , ๐‘ข(๐‘ฅ, 0) = ๐‘ ๐‘–๐‘› ( ๐œ‹๐‘ฅ/2 ) + 3 ๐‘ ๐‘–๐‘› ( 5๐œ‹๐‘ฅ/2 ) , 0 < ๐‘ฅ < 2


1
Expert's answer
2021-11-30T12:26:03-0500

"\begin{aligned}
&\frac{\partial u}{\partial t}=1.71 \frac{\partial^{2} u}{\partial x^{2}} ...(1) \\
& u(x, 0)=\sin \left(\frac{\pi x}{2}\right)+3 \sin \left(\frac{5 \pi x}{2}\right) ,0<x<2
\end{aligned}"

By using variable separation method

"\begin{aligned}
& u(x, t)=X(x) \cdot T(t)...(2) \\
\& & \frac{\partial^{2} u}{\partial x^{2}}=X^{\prime } \cdot T, \quad \frac{\partial u}{\partial t}=X T^{\prime} \\
& \therefore E q^{n}(1) \text { becomes, } \\
& X T^{\prime}=1.71X''T \\
& \frac{T^{\prime}}{1.71 . T}=\frac{X^{\prime \prime}}{X}=K \\
\Rightarrow & \frac{T^{\prime}}{1.71 T}=K , \frac{X^{\prime \prime}}{X}=K \quad(\text { say })
\end{aligned}"

"\begin{aligned}
&\text { Solving, } \frac{T^{\prime}}{1.71 . T}=k\\
&\Rightarrow T^{\prime}=1.71 \mathrm{KT}\\
&\Rightarrow T^{\prime}-1.71 \mathrm{KT}=0\\
&\text { Put } T^{\prime}=m \quad \& \quad D(T)=0\\
&\Rightarrow m-1.71 k=0 \text {. }\\
&\Rightarrow m=1.71 \mathrm{k}\\
&\therefore T(t)=A e^{1.71 k t}...(3)\\
&\text { \& solving } \frac{X^{\prime \prime}}{X}=k\\
&\Rightarrow \quad X^{\prime \prime}-k X=0\\
&\text { Put } X^{\prime}=m \text {. }\\
&\Rightarrow m^{2}-k=0\\
&\Rightarrow m^{2}=k\\
&\Rightarrow m=\pm k\\
&\therefore X(x)=B e^{k x}+C e^{-k x}...(4)
\end{aligned}"

Put eqn (3) \& (4) in eqn (2), we get

"\begin{aligned}
&u(x, t)=A \cdot e^{1,71 k t} \cdot\left(B e^{k x}+c e^{-k x}\right)...(5) \\
&\& B y \text { condition. } \\
&u(x, 0)=\sin \left(\frac{\pi x}{2}\right)+3 \sin \left(\frac{5 \pi x}{2}\right) \quad 0<x<z \\
&\quad \text { i.e. } \quad x=1 \\
&\Rightarrow u(x, 0)=\sin \left(\frac{\pi}{2}\right)+3 \sin \left(\frac{5 \pi}{2}\right) \\
&u(x, 0)=1+3(1) \\
&u(x, 0)=1+3=4 \\
&u(x, 0)=4
\end{aligned}"

"\begin{aligned}
\therefore u(x, 0) &=A e^{1.71 k(0)}\left(B e^{k x}+c e^{-k x}\right)=4 \\
\Rightarrow A(1)\left(B e^{k x}+C e^{-k x}\right) &=4 \\
\Rightarrow A &=\frac{4}{B e^{k x}+c e^{-k x}}
\end{aligned}"

Eqn (5) becomes

ย 

"u(x, t)=\frac{4 e^{1.71 k t}}{B e^{k x}+c e^{k x}} \cdot\left(B e^{k x}+c e^{-k x}\right)"


ย "u(x, t)=4 e^{1.71 k t}"

ย